A body which is uniformly retarded comes to rest in 10 seconds after travelling a distance of 20m. Calculate its initial velocity


1
Expert's answer
2021-10-03T16:17:14-0400

The initial velocity can be found from the following kinematic equation (see https://www.khanacademy.org/science/physics/one-dimensional-motion/kinematic-formulas/a/what-are-the-kinematic-formulas):


d=v+v02td = \dfrac{v + v_0}{2}t

where d=20md = 20m is the distance, v=0m/sv = 0m/s is the final speed, v0v_0 is the initial speed, and t=10st = 10s is the time. Expressing v0v_0, obtain:


v0=2dt=220m10s=4m/sv_0 = \dfrac{2d}{t} = \dfrac{2\cdot 20m}{10s} = 4m/s

Answer. 4m/s.